# Data Reconstruction: When You See It and When You Don’t

### v1

@@ Acknowledgments
Haim Kaplan was partially supported by Israel Science Foundation (grant 1595/19) and the Blavatnik Family Foundation. Yishay Mansour’s work partially funded from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 882396), by the Israel Science Foundation (grant number 993/17), Tel Aviv University Center for AI and Data Science (TAD), and the Yandex Initiative for Machine Learning at Tel Aviv University. Shay Moran is a Robert J. Shillman Fellow; he acknowledges support by ISF grant 1225/20, by BSF grant 2018385, by an Azrieli Faculty Fellowship, by Israel PBC-VATAT, by the Technion Center for Machine Learning and Intelligent Systems (MLIS), and by the European Union (ERC, GENERALIZATION, 101039692). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. Kobbi Nissim’s work partially funded by NSF grant No. 2217678 and by a gift to Georgetown University. Uri Stemmer was Partially supported by the Israel Science Foundation (grant 1871/19) and the Blavatnik Family foundation. Eliad Tsfadia’s work supported in part by a gift to Georgetown University.
Appendix A Differential privacy as Narcissus resiliency
@@
In Appendix E, we demonstrate that differential privacy can be expressed within the framework of Narcissus resiliency (as defined in Definition 1.1.2); however, this translation does incur some loss in terms of the resulting parameters. To overcome this, we introduce a slight variant of Narcissus resiliency that differs from Definition 1.1.2 in that the family of distributions ${\cal F}$ is of distributions over pairs of datasets:
@@
Note that the above theorem refers to the variant of Narcissus resiliency as in Definition A.
@@
Appendix B Computational Narcissus resiliency
@@
We are generally interested in Definition B only under the restrictions that ${\cal M}$ and $R$ are computationally efficient, and every distribution in ${\cal F}$ can be sampled in ${\rm poly}(\lambda)$ time.
B.1 One-wayness as computational Narcissus resiliency
@@
B.2 Ciphertext Indistinguishability (symmetric key) as Narcissus resiliency
@@
B.3 Ciphertext Indistinguishability (public key) as Narcissus resiliency
@@
Recall that EAV-security (Definition B.2) states that, for any two messages $m_{0}$ or $m_{1}$, an adversary that sees a ciphertext $c$ (and nothing else) cannot distinguish between whether this ciphertext encrypts the message $m_{0}$ or $m_{1}$. The following definition extends this to the public-key setting, where the adversary has access to the public key (which effectively gives it access to an encryption oracle for free).
@@
Fix an adversary ${\cal A}=({\cal B},{\cal C})$ as in Definition B.3 and a security parameter $\lambda$. First note that
@@
Appendix C On the composability of Narcissus resiliency

### v2

@@ Acknowledgments
Edith Cohen was partially supported by Israel Science Foundation (grant 1595/19 and 1156/23). Haim Kaplan was partially supported by Israel Science Foundation (grant 1595/19 and 1156/23) and the Blavatnik Family Foundation. Yishay Mansour’s work partially funded from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 882396), by the Israel Science Foundation (grant number 993/17), Tel Aviv University Center for AI and Data Science (TAD), and the Yandex Initiative for Machine Learning at Tel Aviv University. Shay Moran is a Robert J. Shillman Fellow; he acknowledges support by ISF grant 1225/20, by BSF grant 2018385, by an Azrieli Faculty Fellowship, by Israel PBC-VATAT, by the Technion Center for Machine Learning and Intelligent Systems (MLIS), and by the European Union (ERC, GENERALIZATION, 101039692). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. Kobbi Nissim’s work partially funded by NSF grant No. 2217678 and by a gift to Georgetown University. Uri Stemmer was Partially supported by the Israel Science Foundation (grant 1419/24) and the Blavatnik Family foundation. Eliad Tsfadia’s work supported by a gift to Georgetown University.
[…]
Appendix A Example of a Narcissus Resilient Mechanism that is not Differentially Private
Let ${\cal X}$ be a data domain, let ${\cal F}$ be the family of all i.i.d. distributions over ${\cal X}^{n}$, and let $q\colon{\cal X}\rightarrow\{0,1\}$ be a counting query. In the following we describe a deterministic mechanism that is clearly not differentially private but is $(\varepsilon,\delta,{\cal F})$-$R$-Narcissus-resilient for the predicate $R$ that given $S=(x_{1},\ldots,x_{n})$ and $z=(i,b)\in[n]\times\{0,1,\perp\}$ as inputs, outputs $1$ iff $q(x_{i})=b$ (i.e., the adversary’s task is to predict the value of the counting query on one data element, and we allow it to “fail” intentionally by outputing $\perp$ as a prediction). The mechanism is describe below.
[…]
Input: $S=(x_{1},\ldots,x_{n})\in{\cal X}^{n}$. Parameters: $\varepsilon,\delta>0$. Operation: 1. Compute $\tilde{p}=\frac{1}{n}\sum_{i=1}^{n}q(x_{i})$ (the empirical mean). 2. Let $\tau=(\gamma^{2}+2\gamma)\cdot\frac{\ln(4/\delta)}{n}$ for $\gamma=\frac{6e^{\varepsilon}(1-\delta)-2}{(1-\delta)e^{\varepsilon}-1}$ (note that $\gamma=\tilde{O}\mathopen{}\mathclose{{}\left(1/\varepsilon}\right)$ for $\varepsilon<1$ and $\delta\ll 1$). 3. If $\tilde{p}<\tau$ or $\tilde{p}>1-\tau$, output $\perp$ (i.e., abort). 4. Otherwise, output $\tilde{p}$.
Namely, as long as the empirical mean is not too small or too large, the mechanism simply outputs it without any additional noise (that is required for a DP mechanism). We remark that cutting the edges is necessary, as otherwise, the mechanism wouldn’t have been Narcissus-resilient. [^6] Let ${\cal D}={\cal P}^{n}$ and assume $p:={\mathrm{Pr}}_{x\sim{\cal P}}\mathopen{}\mathclose{{}\left[q(x)=1}\right]=1/n^{2}$. Consider the attacker ${\cal A}$ that given $\tilde{p}$ as input, samples $i\leftarrow[n]$ and outputs $(i,1)$ if $\tilde{p}>0$ and otherwise outputs $(i,\perp)$. The probability it reconstructs w.r.t. $S$ is ${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\tilde{p}>0}\right]\cdot\Omega(1/n)=\Omega(1/n^{2})$, but the probability it reconstructs w.r.t. $T$ (fresh i.i.d. samples) is ${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\tilde{p}>0}\right]\cdot 1/n^{2}=O(1/n^{3})$. Yet, we show that by cutting it when it is only $\tilde{O}\mathopen{}\mathclose{{}\left(\frac{1}{\varepsilon^{2}n}}\right)$ close to the edges (for $\varepsilon<1$ and $\delta\ll 1$) ensures $(\varepsilon,\delta,{\cal F})$-$R$-Narcissus-resiliency.
[…]
Let $\varepsilon,\delta>0$ such that $\gamma(\varepsilon,\delta)\geq 6$ and assume that $n\geq 4(\gamma^{2}+2\gamma)\ln(1/\delta)$ (i.e., $\tau\leq 1/4$). Then the mechanism ${\cal M}$ (with parameters $\varepsilon,\delta$) is $(\varepsilon,\delta,{\cal F})$-$R$-Narcissus-resilient.
Proof.
Fix an i.i.d. distribution ${\cal D}={\cal P}^{n}$ and let $p={\mathrm{Pr}}_{x\sim{\cal P}}\mathopen{}\mathclose{{}\left[q(x)=1}\right]$. We assume without loss of generality that $p\leq 1/2$, as the proof for the case $p\geq 1/2$ holds by symmetry.
Define the random variables $\mathbb{S}=(\mathbb{x}_{1},\ldots,\mathbb{x}_{n})\sim{\cal P}^{n}$, $\mathbb{T}=(\mathbb{x}_{1}^{\prime},\ldots,\mathbb{x}_{n}^{\prime})\sim{\cal P}^{n}$, $\tilde{\mathbb{p}}=\frac{1}{n}\sum_{i=1}^{n}q(\mathbb{x}_{i})$ and $\mathbb{y}={\cal M}(\mathbb{S})$ (note that $\mathbb{y}$ is independent of $\mathbb{T}$). We prove the claim by showing that for every $i\in[n]$:
${\mathrm{Pr}}_{y\sim\mathbb{y}}\mathopen{}\mathclose{{}\left[q(\mathbb{x}_{i})|_{\mathbb{y}=y}\>\approx_{\varepsilon,\delta/2}\>Bern(p)}\right]\geq 1-\delta/2,$ (4)
where $Bern(p)$ denotes the Bernoulli distribution that outputs $1$ w.p. $p$. Given Equation 4 and since $q(\mathbb{x}_{i}^{\prime})\equiv Bern(p)$, we deduce that
$(\mathbb{y},q(\mathbb{x}_{i}))\approx_{\varepsilon,\delta}(\mathbb{y},q(\mathbb{x}_{i}^{\prime})).$ (5)
In order to complete the proof of the claim using (5), we now fix an adversary ${\cal A}\colon([0,1]\cup\{\perp\})\rightarrow([n]\times\{0,1,\perp\})$ which can be described by two (randomized) functions $I,B$ that given $y$ as input, it first samples an index $i\sim I(y)\in[n]$ and then a bit $b\sim B(i,y)\in\{0,1,\perp\}$. For $i\in[n]$ we define the random function $F_{i}$ that given $x\in{\cal X}$ and $y\in[0,1]$ as inputs, samples $b\sim B(i,y)$ and outputs $\mathds{1}{\{q(x)=b\}}$. In words, $F_{i}(x,y)$ emulates an execution of ${\cal A}(y)$ conditioned on ${\cal A}$ aiming to predict the $i^{\rm th}$ bit, and outputs $1$ iff it outputs $x$. Next, define the random variables $\mathbb{z}=(\mathbb{i},\mathbb{b})=\mathsf{A}(\mathbb{y})$. Now for every $i\in\operatorname{Supp}(\mathbb{i})$, by the definition of $F_{i}$, it holds that $R(\mathbb{S},\mathbb{z})|_{\mathbb{i}=i}\equiv F_{i}(q(\mathbb{x}_{i}),\mathbb{y})$ and $R(\mathbb{T},\mathbb{z})|_{\mathbb{i}=i}\equiv F_{i}(q(\mathbb{x}_{i}^{\prime}),\mathbb{y})$. Thus, by Equation 5 and post-processing we deduce that $R(\mathbb{S},\mathbb{z})|_{\mathbb{i}=i}\approx_{\varepsilon,\delta}R(\mathbb{T},\mathbb{z})|_{\mathbb{i}=i}$. Since the above holds for every $i\in\operatorname{Supp}(\mathbb{i})$, we conclude that it also holds for $i\sim\mathbb{i}$ which yields that $R(\mathbb{S},\mathbb{z})\approx_{\varepsilon,\delta}R(\mathbb{T},\mathbb{z})$, as required.
It is left to prove Equation 4. Recall that $\tilde{\mathbb{p}}=\frac{1}{n}\sum_{i=1}^{n}q(\mathbb{x}_{i})$ where $\sum_{i=1}^{n}q(\mathbb{x}_{i})$ is distributed as a binomial $B(n,p)$. A standard tail inequality for binomial distributions ([ChungLu], Lemma 2.1) implies that
$\forall t>0:\quad{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|\tilde{\mathbb{p}}-p}\right|>t}\right]\leq 2\cdot e^{-\frac{t^{2}n}{2(p+t/3)}}.$ (6)
In the following, let $\Delta=2\gamma\cdot\frac{\ln(4/\delta)}{n}$ and recall that $\tau=\gamma^{2}\cdot\frac{\ln(4/\delta)}{n}+\Delta$ (Step 2 of Appendix A). We now split into three cases.
The case $p\in[0,\tau-\Delta]$.
By Equation 6 it holds that
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|\tilde{\mathbb{p}}-p}\right|>\Delta}\right]\leq 2\cdot e^{-\frac{\Delta^{2}n}{2(p+\Delta/3)}}\leq 2\cdot e^{-\frac{4\gamma^{2}\ln^{2}(4/\delta)/n}{2(\gamma^{2}\ln(4/\delta)/n+2\gamma\ln(4/\delta)/(3n))}}\leq 2\cdot e^{-\ln(4/\delta)}\leq\delta/2,$
where in the penultimate inequality we used the assumption $\gamma\geq 6$. Therefore, we deduce that ${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathbb{y}=\perp}\right]\geq 1-\delta/2$. The latter implies that ${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[q(\mathbb{x}_{i})=1\mid\mathbb{y}=\perp}\right]\in p\pm\delta/2$ which yields that $q(\mathbb{x}_{i})|_{\mathbb{y}=\perp}\approx_{0,\delta/2}Bern(p)$.
[…]
Let $\lambda_{p}:=2\cdot\sqrt{\frac{p\cdot\ln(4/\delta)}{n}}$ and observe that $\lambda_{\tau-\Delta}=\Delta$ and $\Delta<\lambda_{\tau+\Delta}\leq 2\Delta$. In addition, since $p\geq\tau+2\Delta$ it holds that $\lambda_{p}<p$. Thus Equation 6 implies that
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|\tilde{\mathbb{p}}-p}\right|>\lambda_{p}}\right]\leq 2\cdot e^{-\frac{\lambda_{p}^{2}n}{2(p+\lambda_{p}/3)}}\leq 2\cdot e^{-\frac{4p\cdot\ln(4/\delta)}{4p}}\leq\delta/2.$ (7)
Furthermore, note that for every $\tilde{p}\in p\pm\lambda_{p}$ it holds that
$\max\{\frac{p}{\tilde{p}},\frac{\tilde{p}}{p}\}$ $\leq\frac{\max\{p,\tilde{p}\}}{\min\{p,\tilde{p}\}}\leq\frac{p}{p-\lambda_{p}}\leq\frac{\tau+\Delta}{\tau+\Delta-\lambda_{\tau+\Delta}}\leq\frac{\tau+\Delta}{\tau-\Delta}$
$\leq\frac{\tau-2\Delta}{(1-\delta)(\tau-4\Delta)}=\frac{\gamma-2}{(1-\delta)(\gamma-6)}=\frac{\frac{6e^{\varepsilon}(1-\delta)-2}{(1-\delta)e^{\varepsilon}-1}-2}{(1-\delta)(\frac{6e^{\varepsilon}(1-\delta)-2}{(1-\delta)e^{\varepsilon}-1}-6)}=e^{\varepsilon},$ (8)
where the second, third, and fifth inequalities hold since for every $a>b>c>0$ we have $\frac{a}{b}\leq\frac{a-c}{b-c}$.
We deduce that for every $\tilde{p}\in p\pm\lambda_{p}$ it holds that $q(\mathbb{x}_{i})|_{\mathbb{y}=\tilde{p}}\approx_{\varepsilon,0}Bern(p)$, and thus conclude that
${\mathrm{Pr}}_{y\sim\mathbb{y}}\mathopen{}\mathclose{{}\left[q(\mathbb{x}_{i})|_{\mathbb{y}=y}\>\approx_{\varepsilon,0}\>Bern(p)}\right]$ $\geq(1-\delta/2){\mathrm{Pr}}_{y\sim\mathbb{y}|_{\mathopen{}\mathclose{{}\left|\tilde{\mathbb{p}}-p}\right|\leq\lambda_{p}}}\mathopen{}\mathclose{{}\left[q(\mathbb{x}_{i})|_{\mathbb{y}=y}\>\approx_{\varepsilon,0}\>Bern(p)}\right]$
$=(1-\delta/2){\mathrm{Pr}}_{\tilde{p}\sim\tilde{\mathbb{p}}|_{\mathopen{}\mathclose{{}\left|\tilde{\mathbb{p}}-p}\right|\leq\lambda_{p}}}\mathopen{}\mathclose{{}\left[q(\mathbb{x}_{i})|_{\mathbb{y}=\tilde{p}}\>\approx_{\varepsilon,0}\>Bern(p)}\right]$
$=1-\delta/2,$
where the inequality holds by Equation 7 and the first equality holds since $\mathopen{}\mathclose{{}\left|\tilde{\mathbb{p}}-p}\right|\leq\lambda_{p}$ implies that $\mathbb{y}=\tilde{\mathbb{p}}$ because the function $f(p)=p-\lambda_{p}$ is monotonic increasing and $f(\tau+2\Delta)>0$.
The case $p\in[\tau-\Delta,\tau+2\Delta]$.
Here we use a truncated version of Lemma 2.1 (inequality 2.1) in [ChungLu], which implies that:
 $\forall\>0\leq a\leq b\leq p:\quad{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\tilde{\mathbb{p}}<p-b\mid\tilde{\mathbb{p}}<p-a}\right]\leq 2\cdot e^{-\frac{(b-a)^{2}n}{2p}}.$ (9)
If $p\in[\tau,\tau+2\Delta]$ we deduce by Equation 9 that
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\tilde{\mathbb{p}}<\tau-2\Delta\mid\tilde{\mathbb{p}}<\tau}\right]\leq 2\cdot e^{-\frac{4\Delta^{2}n}{2p}}\leq e^{-\frac{4\Delta^{2}n}{2(\tau+2\Delta)}}=2\cdot e^{-\frac{16\gamma^{2}\ln^{2}(4/\delta)}{2(\gamma^{2}+5\gamma)\ln(4/\delta)}}\leq\delta.$
If $p\in[\tau-\Delta,\tau]$ then by Equation 9 we have
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\tilde{\mathbb{p}}<\tau-2\Delta\mid\tilde{\mathbb{p}}<\tau}\right]=\frac{{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\tilde{\mathbb{p}}<\tau-2\Delta}\right]}{{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\tilde{\mathbb{p}}<\tau}\right]}\leq 2\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\tilde{\mathbb{p}}<\tau-2\Delta}\right]\leq 2\cdot e^{-\frac{\Delta^{2}n}{2\tau}}=2\cdot e^{-\frac{4\gamma^{2}\ln^{2}(4/\delta)}{2(\gamma^{2}+\gamma)\ln(4/\delta)}}\leq\delta.$
Overall, in both cases, we conclude that
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[q(\mathbb{x}_{i})=1\mid\mathbb{y}=\perp}\right]={\mathrm{E}}\mathopen{}\mathclose{{}\left[\tilde{\mathbb{p}}\mid\tilde{\mathbb{p}}<\tau}\right]\geq(1-\delta)(\tau-2\Delta).$
This implies that $q(\mathbb{x}_{i})|_{\mathbb{y}=\perp}\>\approx_{\varepsilon,0}\>Bern(p)$ since
$\frac{p}{{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[q(\mathbb{x}_{i})=1\mid\mathbb{y}=\perp}\right]}\leq\frac{p}{(1-\delta)(\tau-2\Delta)}\leq\frac{\tau-\Delta}{(1-\delta)(\tau-2\Delta)}\leq\frac{\tau-2\Delta}{(1-\delta)(\tau-4\Delta)}=e^{\varepsilon}.$
where the last inequality holds by Appendix A.
Furthermore, for every $\tilde{p}\in p\pm 2\Delta$ we have that $q(\mathbb{x}_{i})|_{\mathbb{y}=\tilde{p}}\>\approx_{\varepsilon,0}\>Bern(p)$ since
$\max\{\frac{p}{\tilde{p}},\frac{\tilde{p}}{p}\}\leq\frac{\max\{p,\tilde{p}\}}{\min\{p,\tilde{p}\}}\leq\frac{p}{p-2\Delta}\leq\frac{\tau-2\Delta}{\tau-4\Delta}\leq e^{\varepsilon}$
where the second and third inequalities hold since for every $a>b>c>0$ we have $\frac{a}{b}\leq\frac{a-c}{b-c}$, and the last one holds by Appendix A.
We thus conclude that
${\mathrm{Pr}}_{y\sim\mathbb{y}}\mathopen{}\mathclose{{}\left[q(\mathbb{x}_{i})|_{\mathbb{y}=y}\>\approx_{\varepsilon,0}\>Bern(p)}\right]$ $\geq{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|p-\tilde{\mathbb{p}}}\right|\leq 2\Delta}\right]\cdot{\mathrm{Pr}}_{y\sim\mathbb{y}|_{\mathopen{}\mathclose{{}\left|p-\tilde{\mathbb{p}}}\right|\leq 2\Delta}}\mathopen{}\mathclose{{}\left[q(\mathbb{x}_{i})|_{\mathbb{y}=y}\>\approx_{\varepsilon,0}\>Bern(p)}\right]$
$={\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|p-\tilde{\mathbb{p}}}\right|\leq 2\Delta}\right]\cdot 1$
$\leq 2\cdot e^{-\frac{4\Delta^{2}n}{2(p+2\Delta/3)}}\leq 2\cdot e^{-\frac{4\Delta^{2}n}{2(\tau+2\Delta+2\Delta/3)}}\leq 2\cdot e^{-\frac{4\gamma^{2}\ln^{2}(4/\delta)}{2(\gamma^{2}+6\gamma)\ln(4/\delta)}}\leq\delta/2,$
as required. The second inequality holds by Equation 6, the third one holds since $p\leq\tau+2\Delta$, and the last one holds since $\gamma\geq 6$.
$\Box$
Appendix B Differential privacy as Narcissus resiliency
@@
In Appendix F, we demonstrate that differential privacy can be expressed within the framework of Narcissus resiliency (as defined in Definition 1.1.2); however, this translation does incur some loss in terms of the resulting parameters. To overcome this, we introduce a slight variant of Narcissus resiliency that differs from Definition 1.1.2 in that the family of distributions ${\cal F}$ is of distributions over pairs of datasets:
@@
Note that the above theorem refers to the variant of Narcissus resiliency as in Definition B.
@@
Appendix C Computational Narcissus resiliency
@@
We are generally interested in Definition C only under the restrictions that ${\cal M}$ and $R$ are computationally efficient, and every distribution in ${\cal F}$ can be sampled in ${\rm poly}(\lambda)$ time.
C.1 One-wayness as computational Narcissus resiliency
@@
C.2 Ciphertext Indistinguishability (symmetric key) as Narcissus resiliency
@@
C.3 Ciphertext Indistinguishability (public key) as Narcissus resiliency
@@
Recall that EAV-security (Definition C.2) states that, for any two messages $m_{0}$ or $m_{1}$, an adversary that sees a ciphertext $c$ (and nothing else) cannot distinguish between whether this ciphertext encrypts the message $m_{0}$ or $m_{1}$. The following definition extends this to the public-key setting, where the adversary has access to the public key (which effectively gives it access to an encryption oracle for free).
@@
Fix an adversary ${\cal A}=({\cal B},{\cal C})$ as in Definition C.3 and a security parameter $\lambda$. First note that
@@
Appendix D On the composability of Narcissus resiliency